# Diffusion Beats Autoregressive in Data-Constrained Settings

### v4

@@ 1 Introduction
    Diffusion models have a much higher effective epoch count. Muennighoff et al. muennighoff2023scaling fit scaling laws for AR models in data-constrainted settings and define $R_{D}^{*}$ as a learned constant that characterizes the number of epochs after which training more epochs results in significantly diminished returns. For autoregressive models, they estimate $R_{D}^{*}\approx 15$ In contrast, we find $R_{D}^{*}\approx 500$ for diffusion models, suggesting they can benefit from repeated data over far more epochs without major degradation (Section 4.2).
@@ 2 Related Work
Diffusion models, originally developed for image generation (ho2020denoising, ), have recently been adapted to text, offering a fundamentally different paradigm for language modeling (austin2021structured, ; li2022diffusion, ; gulrajani2023likelihood, ). Broadly, diffusion language models fall into two categories: continuous and discrete. Continuous approaches (gulrajani2023likelihood, ) inject Gaussian noise in the forward process, whereas discrete methods (austin2021structured, ) corrupt tokens with noise sampled from distributions such as Bernoulli. Among the two classes, continuous diffusion has proven more difficult to scale on language data gulrajani2023likelihood ; lou2023discrete . In contrast, recent advances in discrete diffusion—particularly masked diffusion—have shown encouraging results. Recent work (arriola2025block, ; dieleman2023language, ; sahoo2024simple, ; lou2023discrete, ) has significantly narrowed the performance gap between diffusion and AR models. Notably, LLaDA (nie2025large, ) scales masked diffusion models to 8B parameters and achieves results similar to LLaMA3-8B across both pretraining and instruction-tuned evaluations. Furthermore, Nie et al. (nie2024scaling, ) provide scaling law analysis showing that diffusion models follow similar power-law trends as AR models, though they may require up to 16$\times$ more compute under single-epoch training, Swerdlow et al. (swerdlow2025unidisc, ) find similar trends on multimodal data containing both image and text. However, these evaluations are restricted to single-pass training and do not examine the data-constrained, multi-epoch regimes which is the focus of our work.
@@ 3.2 Modeling Details for AR and Masked Diffusion
[…]
@@ 4.1 Does Diffusion Beat AR in Data-Constrained Settings?
Prior comparisons between diffusion and autoregressive (AR) language models have largely focused on the single-epoch regime, where each token is seen only once during training nie2024scaling ; swerdlow2025unidisc . In this setting, diffusion models are consistently reported to require substantially more training compute ($C\sim 6ND$) than AR models to achieve comparable validation loss. For instance, Nie et al. (nie2024scaling, ) and Swerdlow et al. (swerdlow2025unidisc, ) derive scaling laws showing that masked diffusion models can require up to 16$\times$ more compute than AR counterparts.
@@ 5 Discussion
Why do diffusion models outperform autoregressive (AR) models in low-data regimes?
We hypothesize that the key advantage stems from the random factorizations in diffusion objective, which serves as a form of implicit data augmentation. Unlike AR models, which are trained on a single, fixed left-to-right ordering, masked diffusion models are exposed to a wide variety of token prediction tasks and orderings during training—including, the left-to-right order itself. This broader distribution over prediction tasks encourages better generalization and more effective use of each training example. Analogous to image-based tasks where techniques like random cropping or color jittering boost generalization  caron2021emerging ; cubuk2020randaugment ; he2022masked , we believe the corruption process in diffusion plays a similar role. By learning to denoise across many corruptions, the model extracts richer signal per example over time, resulting in improved data efficiency. We further analyze this question in Section 9 in Appendix.
@@ 6 Limitations
In this work, we examined two extremes of generative modeling: masked diffusion models, which learn random token orderings and are more data-efficient, and autoregressive (AR) models, which follow a fixed left-to-right order and are more compute-efficient. While our results highlight a clear trade-off, this need not be binary—hybrid models that interpolate between AR and diffusion would offer a better balance. Although prior works have explored such hybrids (arriola2025block, ; hoogeboom2021autoregressive, ), they have not been evaluated through the lens of data-compute efficiency. Additionally, our scaling laws are currently fit over a limited range of unique data sizes; extending them to larger regimes may improve predictive accuracy and reveal further insights.
@@ 8 Acknowledgement
We thank Alexander Li for his valuable insights and detailed feedback on the manuscript. We are also grateful to Zheyang Qin for his help in improving the figures. Finally, we thank Niklas Muennighoff, Simo Ryu and Colin Raffel for their helpful comments on the final draft of the paper. We thank Stella Biderman for pointing out the inconsistent unit used in the Critical Compute equation. This work was supported in part by ONR MURI N00014-22-1-2773, ONR N00014-22-1-2096, and AFOSR FA9550-23-1-0747.

### v5

@@ 1 Introduction
    Diffusion models have a much higher effective epoch count. Muennighoff et al. muennighoff2023scaling fit scaling laws for AR models in data-constrainted settings and define $R_{D}^{*}$ as a learned constant that characterizes the number of epochs after which training more epochs results in significantly diminished returns. For autoregressive models, they estimate $R_{D}^{*}\approx 15$ . In contrast, we find $R_{D}^{*}\approx 500$ for diffusion models, suggesting they can benefit from repeated data over far more epochs without major degradation (Section 4.2).
@@ 2 Related Work
[…]
@@ 3.2 Modeling Details for AR and Masked Diffusion
For masked diffusion we first sample a masking ratio $r\sim\mathcal{U}(0,1)$ and construct a corrupted sequence $\tilde{x}$ by independently replacing each token with [MASK] with probability $r$. Let $\mathcal{M}=\{\,i:\tilde{x}_{i}=\textsc{[MASK]}\,\}$ be the set of masked positions. The loss is then
@@ 4.1 Does Diffusion Beat AR in Data-Constrained Settings?
[…]
@@ 4.4 Downstream Results
4.5 Why do Diffusion models outperform AR models in data-constrained settings?
To better understand why diffusion models are more data-efficient than autoregressive (AR) models, we conducted a series of controlled experiments aimed at isolating the core source of diffusion’s advantage.
Our hypothesis is that masked diffusion models benefit from a form of implicit data augmentation arising from their randomized masking and denoising objective. Unlike AR models, which are trained exclusively on a fixed left-to-right factorization, diffusion models are exposed to a diverse set of conditional prediction tasks due to random masking patterns, enabling them to learn to generate tokens in varying orders.
(a) Validation loss under varying attention dropout levels in AR training.
(b) Validation loss under varying token masking levels in AR training.
Figure 7: Impact of common data augmentations on AR models. Despite applying attention dropout and token masking, AR models still overfit and underperform compared to diffusion models. We believe this gap arises because diffusion models learn random factorizations of the joint distribution, rather than a fixed left-to-right ordering.
To test whether this diversity could be replicated in AR models through explicit augmentation, we first experimented with standard perturbation-based techniques during AR training. Specifically, we applied (i) attention dropout, where a random fraction of attention weights was dropped at varying levels (25%, 50%, 75%), and (ii) token masking, where a subset of input tokens was masked by zeroing out their attention weights across all layers, while retaining the next-token prediction objective. As shown in Figure 7(a) and Figure 7(b), neither of these augmentations improved validation loss. In all cases, AR models continued to overfit and remained significantly behind diffusion models trained for longer epochs.
Figure 8: Validation loss improves as the number of token orderings NN increases in AR training. At $N=16$, performance approaches that of diffusion models.
We next investigated whether the key factor behind diffusion’s performance lies in its exposure to diverse token orderings. To evaluate this, we trained AR models with varying numbers of token orderings: $N=1$ corresponds to standard left-to-right training, while $N=k$ includes the left-to-right order along with k1k{-}1 additional random permutations of the token sequence. As shown in Figure 8, increasing $N$ consistently reduced validation loss and delayed overfitting. At $N=16$, the AR models approached the performance of diffusion models, suggesting that training over diverse orderings is indeed a major driver of diffusion’s data efficiency.
These findings support our interpretation that diffusion models outperform AR models in low-data regimes because they are implicitly trained over a rich distribution of conditional prediction tasks. While AR models can partially recover this behavior via explicit random orderings, they remain fundamentally constrained by their fixed prediction structure.
Finally, this analysis suggests a natural interpolation between the two paradigms. By controlling the diversity of prediction tasks—through masking or reordering—it may be possible to design hybrid models that trade off between compute efficiency (AR-like behavior) and data efficiency (diffusion-like behavior). Exploring this continuum offers a promising direction for future research. More details on how we permute the sequence can be found in 11
@@ 6 Limitations
In this work, we examined two extremes of generative modeling: masked diffusion models, which learn over random condition prediction tasks and are more data-efficient, and autoregressive (AR) models, which follow a fixed left-to-right order and are more compute-efficient. While our results highlight a clear trade-off, this need not be binary—hybrid models that interpolate between AR and diffusion would offer a better balance. Although prior works have explored such hybrids (arriola2025block, ; hoogeboom2021autoregressive, ), they have not been evaluated through the lens of data-compute efficiency. We explore part of this question in Section 4.5, however it will be useful to study this in more detail. Additionally, our scaling laws are currently fit over a limited range of unique data sizes; extending them to larger regimes may improve predictive accuracy and reveal further insights.
@@ 8 Acknowledgement
We thank Alexander Li for his valuable insights and detailed feedback on the manuscript. We are also grateful to Zheyang Qin for his help in improving the figures. Finally, we thank Niklas Muennighoff, Simo Ryu and Colin Raffel for their helpful comments on the final draft of the paper. We thank Stella Biderman for pointing out the inconsistent unit used in the Critical Compute equation. We thank Lucas Beyer and You Jiacheng, in suggesting the experiments in Section 4.5, that helped us explain why diffusion models would be more data-efficient. This work was supported in part by ONR MURI N00014-22-1-2773, ONR N00014-22-1-2096, and AFOSR FA9550-23-1-0747.